# references.bib

### v1

@@ 2.3 Dual formulation of the reduced risk
In eqs. 9 and 10, the objectives $\mathcal{L}^{\lambda}_{f}$ and $\mathcal{L}_{f}^{0}$ are expressed as a the value of a minimisation problem over the outer weights $u$. Taking the dual of those minimization problems, $\mathcal{L}_{f}^{\lambda}$ and $\mathcal{L}^{0}_{f}$ can be expressed as the value of a maximisation problem over the dual variable $\alpha\in L^{2}(\rho)$. In contrast with the primal formulation eq. 9, the dual formulation of Proposition 2.1 has the advantage of conveniently expressing $\mathcal{L}^{\lambda}_{f}$ for both $\lambda>0$ and $\lambda=0$ as the value of an optimization problem over the space $L^{2}(\rho)$ which is independent of $\mu$.
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\langle\alpha,Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)},$ (19)
@@
$\forall\alpha\in L^{2}(\rho),\quad\Phi^{\top}\alpha=\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x).$
@@
Consider $\mu\in\mathcal{P}(\Omega)$ and $\lambda>0$. First, by definition of $\Phi^{\top}$ we have for every $\alpha\in L^{2}(\rho)$ and every $u\in L^{1}(\mu)$ that $\int_{\Omega}(\Phi^{\top}\alpha)u\mathrm{d}\mu=\langle\alpha,\Phi_{\mu}\cdot u\rangle_{L^{2}(\rho)}$ i.e. $\Phi^{\top}$ is the adjoint of $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\mu)$. Also, it follows from the assumption on $f$ that the map $\mathcal{I}_{f}:u\in L^{1}(\mu)\mapsto\int_{\Omega}f(u)\mathrm{d}\mu$ is a convex, weakly lower-semicontinuous functional whose Legendre transform is given for $h\in L^{\infty}(\mu)$ by:
@@
$\sup_{\alpha\in L^{2}(\rho)}-\langle\alpha,\Phi_{\mu}\cdot u-Y\rangle_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}=\frac{1}{2\lambda}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}.$
@@
The formula eq. 22 is directly deduced from eq. 19 and the characterization of the RKHS $\mathcal{H}$ in eq. 41. Also eq. 23 is a rewritting of eq. 20 since $\nu^{\lambda}_{f}[\mu]\in\mathcal{M}(\Omega)$ and $h^{\lambda}_{f}[\mu]\in\mathcal{H}$ are related to $u^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ and $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ by $\mathrm{d}\nu^{\lambda}_{f}[\mu]=u^{\lambda}_{f}[\mu]\mathrm{d}\mu\quad\text{and}\quad h^{\lambda}_{f}[\mu]=\Phi^{\top}\alpha^{\lambda}_{f}[\mu]$. ∎
@@ 2.4 Kernel learning in the case of quadratic regularization
For the $L^{2}$-regularization $f(t)=|t|^{2}$, the reduced risk $\mathcal{L}^{\lambda}_{2}[\mu]$ is the value of the ridge regression problem in eq. 4 and for $\lambda>0$, the optimizer is given by $u^{\lambda}_{2}[\mu]=(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$ where $\Phi_{\mu}^{\top}:L^{2}(\rho)\to L^{2}(\mu)$ is the adjoint of the operator $\Phi_{\mu}$ restricted to $L^{2}(\mu)$. Also, the dual problem reads:
$\mathcal{L}^{\lambda}_{2}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\frac{1}{2}\langle\alpha,(K_{\mu}+2\lambda)\alpha\rangle_{L^{2}(\rho)}+\langle\alpha,Y\rangle_{L^{2}(\rho)},$
@@
$\mathcal{L}^{\lambda}_{2}(\mu)=\frac{1}{2}\langle Y,(K_{\mu}+2\lambda)^{-1}Y\rangle_{L^{2}(\rho)}.$ (24)
This is the optimal value of the kernel ridge regression problem with kernel $K_{\mu}$ parameterized by the feature distribution $\mu$. Moreover this parameterization is linear w.r.t. $\mu\in\mathcal{P}(\Omega)$ since considering for $\omega\in\Omega$ the rank the rank-one self-adjoint operator $k(\omega)\coloneqq\phi(\omega,.)\otimes\phi(\omega,.)$ we have:
@@
Therefore, minimizing the reduced risk $\mathcal{L}^{\lambda}_{2}$ over the feature distribution $\mu$ amounts to find the best kernel for solving the ridge regression problem in eq. 4 among convex combinations of “simple” basis kernels $(k(\omega))_{\omega\in\Omega}$ i.e. a Multiple Kernel Learning task. Other convex optimization strategies for solving such task have been studied in [lanckriet2004learning, bach2004multiple].
@@ 3.1 Existence and uniqueness of minimizers
We first investigate existence and uniqueness of minimizers of $\mathcal{L}^{\lambda}_{f}$. Importantly, we use here that $\mathcal{L}^{\lambda}_{f}$ is obtained as a the result of a partial minimization. Namely, for $\lambda\geq 0$ and $\mu\in\mathcal{P}(\Omega)$, we have from eq. 16 that $\mathcal{L}^{\lambda}_{f}(\mu)=\min_{\nu\in\mathcal{M}(\Omega)}\mathcal{E}^{\lambda}_{f}(\nu,\mu)$, where $\mathcal{E}^{\lambda}_{f}$ is defined for $\nu\in\mathcal{M}(\Omega)$ and $\mu\in\mathcal{P}(\Omega)$ by:
$\mathcal{E}^{\lambda}_{f}(\nu,\mu)\coloneqq\left\{\begin{array}[]{cc}\mathrm{D}_{f}(\nu|\mu)+\frac{1}{2\lambda}\|\Phi\star\nu-Y\|^{2}&\text{if $\lambda>0$,}\\ \mathrm{D}_{f}(\nu|\mu)+\iota_{\Phi\star\nu=Y}&\text{if $\lambda=0$.}\end{array}\right.$ (27)
In particular, it follows from variational formulations of the $f$-divergences that $\mathrm{D}_{f}$ is (jointly) convex and lower-semicontinuous w.r.t. its arguments $(\nu,\mu)\in\mathcal{M}(\Omega)\times\mathcal{P}(\Omega)$ [rockafellar1971integrals, Thm. 4]. The following Lemma 3.1 uses this fact to establish convexity and lower-semicontinuity of $\mathcal{L}^{\lambda}_{f}$, implying the existence of minimizers. We then discuss cases in which $\mathcal{L}^{\lambda}_{f}$ has in fact a unique minimizer.
@@
Assume $f$ satisfies Assumption 2. Then, for $\lambda\geq 0$, $\mathcal{L}^{\lambda}_{f}:\mathcal{P}(\Omega)\to\mathbb{R}$ is a convex, lower semi-continuous function (w.r.t. the narrow convergence on $\mathcal{M}(\Omega)$).
@@
Thus $\mathrm{D}_{f}$ is a (jointly) convex and lower-semicontinuous function as a supremum of (jointly) convex and lower semi-continuous functions. As a consequence, for $\lambda\geq 0$, $\mathcal{E}^{\lambda}_{f}$ is also (jointly) convex and lower-semicontinuous and the convexity of $\mathcal{L}^{\lambda}_{f}=\min_{\nu}\mathcal{E}^{\lambda}_{f}(\nu,.)$ follows as partial minimization preserves convexity. Also if $(\mu_{n})_{n\geq 0}$ is a sequence in $\mathcal{P}(\Omega)$ converging narrowly to some $\mu\in\mathcal{P}(\Omega)$, then we have $\mathcal{L}^{\lambda}_{f}(\mu_{n})=\mathcal{E}^{\lambda}_{f}(\nu_{n},\mu_{n})$ for some $\nu_{n}\in\mathcal{M}(\Omega)$. Without loss of generality one can assume $\mathcal{L}^{\lambda}_{f}(\mu_{n})$ is bounded, thus $\mathrm{D}_{f}(\nu_{n}|\mu_{n})$ and then $\|\nu_{n}\|_{\mathrm{TV}}$ are bounded as $f$ is superlinear. Then, up to extraction of a subsequence, $(\nu_{n})$ converges narrowly to $\nu\in\mathcal{M}(\Omega)$ and we get by lower semi-continuity of $\mathcal{E}^{\lambda}_{f}$:
@@
which shows that $\mathcal{L}^{\lambda}_{f}$ is lower-semicontinuous.
The above result implies the existence of minimizers of the reduced risk $\mathcal{L}^{\lambda}_{f}$ for every $\lambda\geq 0$ but it does not establish uniqueness and $\mathcal{L}^{\lambda}_{f}$ may, a priori, have several minimizers. However, there are cases in which uniqueness can be ensured, we give two examples:
  • In the teacher student setup where Assumption 1 holds, if $\bar{\nu}$ is a positive measure with $\bar{m}=\bar{\nu}(\Omega)>0$ and if $f$ is nonnegative, strictly convex and such that $f(\bar{m})=0$ then the teacher feature distribution $\bar{\mu}=\bar{\nu}/\bar{m}$ is the unique minimizer of $\mathcal{L}^{\lambda}_{f}$, whatever $\lambda\geq 0$. Indeed, from eq. 18 we have that $\mathcal{L}^{\lambda}_{f}(\bar{\mu})=0$ and $\mathcal{L}^{\lambda}_{f}(\mu)>0$ for every $\mu\neq\bar{\mu}$. With these assumptions, we prove in Theorem 5.1 that the gradient flow of $\mathcal{L}^{\lambda}_{f}$ converges towards the teacher feature distribution $\bar{\mu}$ with an algebraic convergence rate.
@@
We use arguments similar to the one of [wang2024mean, Prop. 3.3]. By duality $\inf_{\mu\in\mathcal{P}(\Omega)}\mathcal{L}^{\lambda}_{r}(\mu)=\inf_{\nu\in\mathcal{M}(\Omega)}\mathcal{G}^{\lambda}_{r}(\nu)$ where $\mathcal{G}^{\lambda}_{r}$ is defined for $\forall\nu\in\mathcal{M}(\Omega)$ by:
@@
The optimality condition gives that $\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}$ is constant and the minimum is attained for $\mu=|\nu|/\|\nu\|_{\mathrm{TV}}$, giving $\int_{\Omega}\left|\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu=\|\nu\|^{r}_{\mathrm{TV}}$. Finally, the map $\nu\mapsto\|\nu\|^{r}_{\mathrm{TV}}$ is strictly convex so that $\mathcal{G}^{\lambda}_{r}$ admits a unique minimizer $\bar{\nu}^{\lambda}_{r}\in\mathcal{M}(\Omega)$, thus $\mathcal{L}^{\lambda}_{r}$ has also a unique minimizer $\bar{\mu}^{\lambda}_{r}\in\mathcal{P}(\Omega)$ and we have the duality relation $\bar{\mu}^{\lambda}_{r}=|\bar{\nu}^{\lambda}_{r}|/\|\bar{\nu}^{\lambda}_{r}\|_{\mathrm{TV}}$. Notably, in the case where Assumption 1 holds and $\lambda=0$, for every $r>1$, we have $\bar{\nu}^{0}_{r}=\bar{\nu}$ and $\bar{\mu}^{0}_{r}=\bar{\mu}$. ∎

### v2

@@ 2.3 Dual formulation of the reduced risk
In Eqs. 17 and 18, the objectives $\mathcal{L}^{\lambda}_{f}$ and $\mathcal{L}_{f}^{0}$ are expressed as the value of a minimization problem over the outer weights $u$. Taking the dual of those minimization problems, $\mathcal{L}_{f}^{\lambda}$ and $\mathcal{L}^{0}_{f}$ can be expressed as the value of a maximization problem over the dual variable $\alpha\in L^{2}(\rho)$. In contrast with the primal formulation Eq. 17, the dual formulation of Proposition 2.1 has the advantage of conveniently expressing $\mathcal{L}^{\lambda}_{f}$ for both $\lambda>0$ and $\lambda=0$ as the value of an optimization problem over the space $L^{2}(\rho)$ which is independent of $\mu$.
@@
$\mathcal{L}^{\lambda}_{f}(\mu)=\max_{\alpha\in L^{2}(\rho)}-\int_{\Omega}f^{*}(\Phi^{\top}\alpha)\mathrm{d}\mu+\left<\alpha,Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)},$ (27)
@@
$\forall\alpha\in L^{2}(\rho),\quad\Phi^{\top}\alpha\coloneqq\int_{\mathbb{R}^{d}}\phi(.,x)\alpha(x)\mathrm{d}\rho(x).$
@@
Consider $\mu\in\mathcal{P}(\Omega)$ and $\lambda>0$. First, by definition of $\Phi^{\top}$ we have for every $\alpha\in L^{2}(\rho)$ and every $u\in L^{1}(\mu)$ that $\int_{\Omega}(\Phi^{\top}\alpha)u\mathrm{d}\mu=\left<\alpha,\Phi_{\mu}\cdot u\right>_{L^{2}(\rho)}$ i.e. $\Phi^{\top}$ is the adjoint of $\Phi_{\mu}:L^{1}(\mu)\to L^{2}(\mu)$. Also, it follows from the assumption on $f$ that the map $\mathcal{I}_{f}:u\in L^{1}(\mu)\mapsto\int_{\Omega}f(u)\mathrm{d}\mu$ is a convex, weakly lower semicontinuous functional whose Legendre transform is given for $h\in L^{\infty}(\mu)$ by:
@@
$\sup_{\alpha\in L^{2}(\rho)}-\left<\alpha,\Phi_{\mu}\cdot u-Y\right>_{L^{2}(\rho)}-\frac{\lambda}{2}\|\alpha\|^{2}_{L^{2}(\rho)}=\frac{1}{2\lambda}\left\|\Phi_{\mu}\cdot u-Y\right\|^{2}_{L^{2}(\rho)}.$
@@
The formula Eq. 30 is directly deduced from Eq. 27 and the characterization of the RKHS $\mathcal{H}$ in Eq. 61. Also Eq. 31 is a rewritting of Eq. 28 since $\nu^{\lambda}_{f}[\mu]\in\mathcal{M}(\Omega)$ and $h^{\lambda}_{f}[\mu]\in\mathcal{H}$ are related to $u^{\lambda}_{f}[\mu]\in L^{1}(\mu)$ and $\alpha^{\lambda}_{f}[\mu]\in L^{2}(\rho)$ by $\mathrm{d}\nu^{\lambda}_{f}[\mu]=u^{\lambda}_{f}[\mu]\mathrm{d}\mu$ and $h^{\lambda}_{f}[\mu]=\Phi^{\top}\alpha^{\lambda}_{f}[\mu]$. ∎
@@ 2.4 Kernel learning in the case of quadratic regularization
For the $L^{2}$-regularization $f(t)=|t|^{2}$, the reduced risk $\mathcal{L}^{\lambda}_{2}[\mu]$ is the value of the ridge regression problem in Eq. 12 and for $\lambda>0$, the optimizer is given by $u^{\lambda}_{2}[\mu]=(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$ where $\Phi_{\mu}^{\top}:L^{2}(\rho)\to L^{2}(\mu)$ is the adjoint of the operator $\Phi_{\mu}$ restricted to $L^{2}(\mu)$. Also, the dual problem in Eq. 27 here reads:
$\mathcal{L}^{\lambda}_{2}(\mu)=\sup_{\alpha\in L^{2}(\rho)}-\frac{1}{2}\left<\alpha,(K_{\mu}+2\lambda)\alpha\right>_{L^{2}(\rho)}+\left<\alpha,Y\right>_{L^{2}(\rho)},$
@@
$\mathcal{L}^{\lambda}_{2}(\mu)=\frac{1}{2}\left<Y,(K_{\mu}+2\lambda)^{-1}Y\right>_{L^{2}(\rho)}.$ (32)
This is the optimal value of the kernel ridge regression problem with kernel $K_{\mu}$, where $K_{\mu}$ is parameterized by the feature distribution $\mu$. Moreover this parameterization is linear in $\mu\in\mathcal{P}(\Omega)$ since considering for $\omega\in\Omega$ the rank-one self-adjoint operator $k(\omega)\coloneqq\phi(\omega,.)\otimes\phi(\omega,.)$ we have:
@@
Therefore, minimizing the reduced risk $\mathcal{L}^{\lambda}_{2}$ over the feature distribution $\mu$ amounts to finding the best kernel for solving the ridge regression problem in Eq. 12 among convex combinations of “simple” basis kernels $(k(\omega))_{\omega\in\Omega}$ i.e. a Multiple Kernel Learning task. Other convex optimization strategies for solving such task have been studied in [lanckriet2004learning, bach2004multiple].
@@ 3.1 Existence and uniqueness of minimizers
We first investigate existence and uniqueness of minimizers of $\mathcal{L}^{\lambda}_{f}$. Importantly, we use here that $\mathcal{L}^{\lambda}_{f}$ is obtained as the result of a partial minimization. Namely, for $\lambda\geq 0$ and $\mu\in\mathcal{P}(\Omega)$, we have from Eq. 24 that $\mathcal{L}^{\lambda}_{f}(\mu)=\min_{\nu\in\mathcal{M}(\Omega)}\mathcal{E}^{\lambda}_{f}(\nu,\mu)$, where $\mathcal{E}^{\lambda}_{f}$ is defined for $\nu\in\mathcal{M}(\Omega)$ and $\mu\in\mathcal{P}(\Omega)$ by:
$\mathcal{E}^{\lambda}_{f}(\nu,\mu)\coloneqq\left\{\begin{array}[]{cc}\mathrm{D}_{f}(\nu|\mu)+\frac{1}{2\lambda}\|\Phi\star\nu-Y\|^{2}&\text{if $\lambda>0$,}\\[10.0pt] \mathrm{D}_{f}(\nu|\mu)+\iota_{\Phi\star\nu=Y}&\text{if $\lambda=0$.}\end{array}\right.$ (35)
In particular, it follows from variational formulations of $f$-divergences that $\mathrm{D}_{f}$ is (jointly) convex and lower semicontinuous w.r.t. its arguments $(\nu,\mu)\in\mathcal{M}(\Omega)\times\mathcal{P}(\Omega)$ [rockafellar1971integrals, Thm. 4]. The following Lemma 3.1 uses this fact to establish convexity and lower semicontinuity of $\mathcal{L}^{\lambda}_{f}$, implying the existence of minimizers. We then discuss cases in which $\mathcal{L}^{\lambda}_{f}$ has in fact a unique minimizer.
@@
Assume $f$ satisfies Assumption 2. Then, for $\lambda\geq 0$, $\mathcal{L}^{\lambda}_{f}:\mathcal{P}(\Omega)\to\mathbb{R}$ is a convex, lower semicontinuous function (w.r.t. the narrow convergence on $\mathcal{M}(\Omega)$).
@@
Thus $\mathrm{D}_{f}$ is a (jointly) convex and lower semicontinuous function as a supremum of (jointly) convex and lower semicontinuous functions. As a consequence, for $\lambda\geq 0$, $\mathcal{E}^{\lambda}_{f}$ is also (jointly) convex and lower semicontinuous. The convexity of $\mathcal{L}^{\lambda}_{f}=\min_{\nu}\mathcal{E}^{\lambda}_{f}(\nu,.)$ follows as partial minimization preserves convexity. Also, if $(\mu_{n})_{n\geq 0}$ is a sequence in $\mathcal{P}(\Omega)$ converging narrowly to some $\mu\in\mathcal{P}(\Omega)$, then we have $\mathcal{L}^{\lambda}_{f}(\mu_{n})=\mathcal{E}^{\lambda}_{f}(\nu_{n},\mu_{n})$ for some $\nu_{n}\in\mathcal{M}(\Omega)$. Without loss of generality one can assume $\mathcal{L}^{\lambda}_{f}(\mu_{n})$ is bounded, thus $\mathrm{D}_{f}(\nu_{n}|\mu_{n})$ and then $\|\nu_{n}\|_{\mathrm{TV}}$ are bounded as $f$ is superlinear. Then, up to extraction of a subsequence, $(\nu_{n})$ converges narrowly to $\nu\in\mathcal{M}(\Omega)$ and we get by lower semicontinuity of $\mathcal{E}^{\lambda}_{f}$:
@@
which shows that $\mathcal{L}^{\lambda}_{f}$ is lower semicontinuous. ∎
The above result implies the existence of minimizers of the reduced risk $\mathcal{L}^{\lambda}_{f}$ for every $\lambda\geq 0$ but it does not establish uniqueness and $\mathcal{L}^{\lambda}_{f}$ may, a priori, have several minimizers. However, there are cases in which uniqueness can be ensured. We give two examples:
  • In the teacher-student setup where Assumption 1 holds, if $\bar{\nu}$ is a positive measure with $\bar{m}=\bar{\nu}(\Omega)>0$ and if $f$ is nonnegative, strictly convex and such that $f(\bar{m})=0$ then the teacher feature distribution $\bar{\mu}=\bar{\nu}/\bar{m}$ is the unique minimizer of $\mathcal{L}^{\lambda}_{f}$, whatever $\lambda\geq 0$. Indeed, from Eq. 26 we have that $\mathcal{L}^{\lambda}_{f}(\bar{\mu})=0$ and $\mathcal{L}^{\lambda}_{f}(\mu)>0$ for every $\mu\neq\bar{\mu}$. With these assumptions, we prove in Theorem 4 that the gradient flow of $\mathcal{L}^{\lambda}_{f}$ converges towards the teacher feature distribution $\bar{\mu}$ with an algebraic convergence rate.
@@
We use arguments similar to the one of [wang2024mean, Prop. 3.3]. By duality $\inf_{\mu\in\mathcal{P}(\Omega)}\mathcal{L}^{\lambda}_{r}(\mu)=\inf_{\nu\in\mathcal{M}(\Omega)}\mathcal{G}^{\lambda}_{r}(\nu)$ where $\mathcal{G}^{\lambda}_{r}$ is defined for $\nu\in\mathcal{M}(\Omega)$ by:
@@
The optimality condition gives that $\frac{\mathrm{d}\mu}{\mathrm{d}|\nu|}$ is constant and the minimum is attained for $\mu=|\nu|/\|\nu\|_{\mathrm{TV}}$, giving $\int_{\Omega}\left|\frac{\mathrm{d}\nu}{\mathrm{d}\mu}\right|^{r}\mathrm{d}\mu=\|\nu\|^{r}_{\mathrm{TV}}$. Finally, the map $\nu\mapsto\|\nu\|^{r}_{\mathrm{TV}}$ is strictly convex so that $\mathcal{G}^{\lambda}_{r}$ admits a unique minimizer $\bar{\nu}^{\lambda}_{r}\in\mathcal{M}(\Omega)$, thus $\mathcal{L}^{\lambda}_{r}$ has also a unique minimizer $\bar{\mu}^{\lambda}_{r}\in\mathcal{P}(\Omega)$ and we have the duality relation $\bar{\mu}^{\lambda}_{r}=|\bar{\nu}^{\lambda}_{r}|/\|\bar{\nu}^{\lambda}_{r}\|_{\mathrm{TV}}$. Notably, in the case where Assumption 1 holds and $\lambda=0$, we have $\bar{\nu}^{0}_{r}=\bar{\nu}$ and $\bar{\mu}^{0}_{r}=\bar{\mu}$ for every $r>1$. ∎